# Dynamic Three-dimensional Simulation of Surface Charging on Rotating Asteroids

### v1

@@ Abstract
Surface charging phenomenon of asteroids, mainly resulting from solar wind plasma and solar radiation, has been studied extensively. However, the influence of asteroid’s rotation on surface charging has yet to be fully understood. Here neural network is established to replace numerical integration, improving the efficiency of dynamic three-dimensional simulation. We implement simulation of rotating asteroids and surrounding plasma environment under different conditions, including quiet solar wind and solar storms, various minerals on asteroid’s surface also be considered. Results show that under typical solar wind, the maximum and minimum potential of asteroids will gradually decrease with their increasing periods, especially when solar wind is obliquely incident. For asteroid has period longer than one week, this decreasing trend will become extremely slow. During solar storm passing, solar wind plasma changes sharply, the susceptibility of asteroid’s surface potential to rotation is greatly pronounced. Minerals on surface also count, plagioclase is the most sensitive mineral among those we explored, while ilmenite seems indifferent to changes in rotation periods. Understanding the surface charging of asteroid under various rotation periods or angles, is crucial for further research into solar wind plasma and asteroid’s surface dust motion, providing a reference for safe landing exploration of asteroids.
@@ 1 Introduction
Surface charging of asteroids belongs to the field of interaction between solar wind and asteroids. Researches in this direction can be roughly divided into three categories: electrification properties and electrostatic driving dynamics of surface dust particles, overall charging properties of asteroid surfaces, and deep charging of asteroids induced by surface charging. Existing researches center on the charged properties (Zimmerman et al., 2016; Oudayer et al., 2019) and floating motion (Hartzell & Scheeres, 2012) of regolith grains on the surface, mostly using magnetohydrodynamic (MHD) or hybrid particle-in-cell (Hybrid PIC) model. Here we focus on the second category. At the terminator between night and day, the maximum electric field strength appears, almost 10 times that of dayside (Xie et al., 2023), in this region, electron temperature is the most important parameter (Stubbs et al., 2014). On the dayside of asteroid, photoelectron current usually dominates, surface potential is always positive, theoretical possibilities have been discussed (Kureshi et al., 2020). Correspondingly, on the night side, charging current mainly comes from the incidence of electrons and ions in solar wind plasma, the surface presents a negative potential, during solar energetic particle (SEP) events, this negative potential may reach thousands (Halekas et al., 2007). However, if enough secondary electrons are emitted, potential of night side can sometimes become positive.
While the theoretical foundation of surface charging is relatively solid, static simulations of asteroids at specific angles are also feasible, how to dynamically model and verify the surface potential of asteroids with rotation remains an urgent problem to be solved. Operating traditional modeling techniques for three-dimensional simulation means that after asteroid rotates a certain angle, surrounding environment and surface potential of the entire asteroid need to be recalculated through integration and solving ordinary differential equations, resulting in a significant rise in computational complexity, analysis of potential under different rotation periods will be extremely challenging.
@@ 3.1 Neural Network
BP neural network is a multi-layer feedforward network trained by error inverse propagation algorithm (Hu & Changliang, 2018). Its topology, as shown in Figure 1, can be divided into three levels: input layer, hidden layer, and output layer. Take mean square error (MSE) as standard, by learning the input-output dataset with steepest gradient descent method, through reverse propagation to continuously adjust weights of each layer, in an effort to minimize errors, it can structure and store massive mapping relationships hidden under a big dataset, owning excellent generalization ability and strong fault tolerance (Liu et al., 2022).
@@
Before training, data normalization is considered, eliminate differences in order of magnitude between parameters, thus avoid unnecessary network errors. ReLU is used as activation function, it can compensate for the problem of Vanishing Gradient that Sigmoid and Tanh have, helping network learn complex patterns in data.
@@ 3.2 Dynamical Multi-scale Model
We use the shape of 2016HO3 with a width-to-length ratio smaller than 0.48 to represent asteroid, regard it as a triaxial ellipsoid with semi-major axis of 29.5, 14, 14m (Li & Scheeres, 2021). As indicated in Figure 2, the simulation domain is a sphere with a radius of 500m, asteroid locates at the center of it, the Z-axis is antiparallel to the spin axis of asteroid.
@@
This process is illustrated in Figure 3.
Operate parameters of PIC simulation (Xie et al., 2023) to verify our method, where eV, secondary electron yield has a peak yield of at the energy of eV. solar wind parameters are at their typical values, with number density of , a velocity of 400km/s, time step is 1.0e-6s, while the simulation period is 0.001s, the angle between the major axis of asteroid and solar wind is fixed. Results are shown in Figure 4.
We can see a positive potential up to 13.6V on the dayside and a negative potential as low as -37.4V on the nightside when the angle , while surface potential is +12V near the subsolar point, and -35V on the nightside with PIC simulation. It can be seen that our method is reliable.
Then we will conduct dynamic simulations asteroids about surface charging under different conditions, the three main categories are shown as follows:
(1) Surface Charging Properties of Rotating Asteroids under Normal Condition
@@
(2) Impact of Rotation on Asteroids during the Passage of Solar Storm
@@
(3) Differences between Asteroids with Various Surface Minerals
| Plagioclase | 5.58 | 238 | 2.8 | 1000 |
| Orthopyroxene | 5.14 | 259 | 2.1 | 700 |
| Ilmenite | 4.29 | 297 | 2.5 | 600 |
Solar wind parameters take that same as (1), rotation periods of asteroid are 1 hour, 6 hours and 1 day.
4 Simulation Results
(1) Surface Charging Properties of Rotating Asteroids under Normal Condition
Figure 5 displays the surface potential of asteroid with three classical attitudes, as well as the distribution of electron density and ion density around it, the rotation period is always 1 day. Consistent with the results of static simulation, potential on the nightside is the lowest when compared with other two conditions. Meanwhile, if the main axis of asteroid is perpendicular to the direction of solar wind incidence, potential on the dayside is obviously higher than other situations.
Apart from period, the direction and angle of rotation can also affect charging results, for example, and are the same situation for asteroids when we use static simulation, but considering that they are rotating, their results differ significantly. Besides, from Figure 6, we can conclude that no matter how long the period is, surface potential tends to be low when solar wind is obliquely incident, this phenomenon will become more obvious as rotation period increases. That’s because when asteroid rotates to , the subsolar point , where often has the maximum potential, is close to the terminator . According to Poisson’s Equation, the sudden change in potential over a short distance will strengthen electric field near asteroid, accelerate electrons to 2.21e7m/s, the resulting higher electron density and electron temperature therefore reduce surface potential. This process continues until electric field and plasma reach a dynamic balance, which will be broken and reshaped after asteroid rotates through a large enough angle.
(2) Impact of Rotation on Asteroids during the Passage of Solar Storm
@@ 4 Simulation Results
Figures 7-8 show results from a simulation with the occurrence of CME. Five attitudes are selected for display, in regard to typical solar wind, the dayside always charged positively, its potential depends on solar elevation angle, will decrease apparently during the process from subsolar point to the terminator, until it turns into a small negative potential on the nightside. When the forward shock exists, solar wind number density and temperature experience a sharp increase, solar wind starts to accelerate. At this point, a sudden drop can be seen in surface potential, photoelectron current is inferior to the current generated by high-energy electrons, the dayside is charged negatively, such as the potential of -52V when , the minimum potential on the nightside can reach -84.54V, close to 20 times that under normal circumstance, at this time, electron density here is . Therefore, even though at the same angle, potential difference between parts of asteroid’s surface can exceed 40V, contributing to exponential growth of electric field strength, as seen in Figure 8, electric field strength near asteroid, particularly the terminator, where there is a sharp transition between positive and negative potentials, has reached the level of 10V/m, greatly enhancing the sensitivity of electrons and ions to the charged properties of asteroid, plasma is extremely unstable, thus the surface potential curve performs strong volatility. It only takes 1/12 period to make the average electron density increase from to , bring about a potential drop of more than 10V.
@@
In addition to magnitude, the distribution of potential is also worth investigating. According to equation (11-12) and kinetic theory of plasma, both electrons and ions during shock achieve larger diffusion coefficients, the effect of drift-diffusion can partially overpower the flow of solar wind, increasing the density around asteroid’s backside, then decrease surface potential there, allowing electrons to gather at the terminator, where has the strongest electric field. In contrast to the virtually uniform negative potential on the nightside during typical solar wind and early CME, potential during shock has a distinct gradient. In most cases, potential at the terminator is clearly lower than other parts, except , when solar elevation angle declines most slowly on the dayside, reducing the potential difference near the terminator, thereby weakening the electric field. Because of the fast-flowing solar wind, ions produce a considerable wake near asteroid’s tail, whereas high-temperature electrons do not. As a result, electron current far exceeds ion current, resulting in the lowest potential here.
@@
Then we change the rotation period. Charging results of asteroids with varying periods under typical solar wind have been discussed in (1). We now focus on the remaining three conditions. Apart from 0.467 hour, we select three additional periods for each stage based on their respective characteristics, results are shown in Table 4, including the maximum potential , the minimum potential , the maximum electric field strength , and the range of potential differences occur on asteroids.
@@
(3) Differences between Asteroids with Various Surface Minerals

### v2

@@ Abstract
Surface charging phenomenon of asteroids, mainly resulting from solar wind plasma and solar radiation, has been studied extensively. However, the influence of asteroid’s rotation on surface charging has yet to be fully understood. Here neural network is established to replace numerical integration, improving the efficiency of dynamic three-dimensional simulation. We implement simulation of rotating asteroids and surrounding plasma environment under different conditions, including quiet solar wind and solar storms, various minerals on asteroid’s surface also be considered. For asteroids with rotation periods comparable to orbital period, effect of orbital motion and obliquity also be studied. Results show that under typical solar wind, the maximum and minimum potential of asteroids will gradually decrease with their increasing periods, especially when solar wind is obliquely incident. For asteroid has period longer than one week, this decreasing trend will become extremely slow. During solar storm passing, solar wind plasma changes sharply, the susceptibility of asteroid’s surface potential to rotation is greatly pronounced. Minerals on surface also count, plagioclase is the most sensitive mineral among those we explored, while ilmenite seems indifferent to changes in rotation periods. Understanding the surface charging of asteroid under various rotation periods or angles, is crucial for further research into solar wind plasma and asteroid’s surface dust motion, providing a reference for safe landing exploration of asteroids.
@@ 1 Introduction
Surface charging of asteroids belongs to the field of interaction between solar wind and asteroids. Researches in this direction can be roughly divided into three categories: electrification properties and electrostatic driving dynamics of surface dust particles, overall charging properties of asteroid surfaces, and deep charging of asteroids induced by surface charging. Existing researches center on the charged properties (Zimmerman et al., 2016; Oudayer et al., 2019) and floating motion (Hartzell & Scheeres, 2013) of regolith grains on the surface, mostly using magnetohydrodynamic (MHD) or hybrid particle-in-cell (Hybrid PIC) model. Here we focus on the second category. At the terminator between night and day, the maximum electric field strength appears, almost 10 times that of dayside (Xie et al., 2023), in this region, electron temperature is the most important parameter (Stubbs et al., 2014). On the dayside of asteroid, photoelectron current usually dominates, surface potential is always positive, theoretical possibilities have been discussed (Kureshi et al., 2020). Correspondingly, on the night side, charging current mainly comes from the incidence of electrons and ions in solar wind plasma, the surface presents a negative potential, during solar energetic particle (SEP) events, this negative potential may reach thousands (Halekas et al., 2007). However, if enough secondary electrons are emitted, potential of night side can sometimes become positive (Halekas et al., 2009).
While the theoretical foundation of surface charging is relatively solid, static simulations of asteroids at specific angles are also feasible, a new grid-free 2D plasma simulation code has been invented (Zimmerman et al., 2014), make it able to discuss interaction between solar wind and a small, irregularly shaped asteroid. But how to dynamically model and verify the surface potential of asteroids with rotation remains an urgent problem to be solved. Operating traditional modeling techniques for three-dimensional simulation means that after asteroid rotates a certain angle, surrounding environment and surface potential of the entire asteroid need to be recalculated through integration and solving ordinary differential equations, resulting in a significant rise in computational complexity, analysis of potential under different rotation periods will be extremely challenging.
@@ 3.1 Neural Network
BP neural network is a multi-layer feedforward network trained by error inverse propagation algorithm (Zhu et al., 2023). Its topology, as shown in Figure 1, can be divided into three levels: input layer, hidden layer, and output layer (Adil et al., 2022). Take mean square error (MSE) as standard, by learning the input-output dataset with steepest gradient descent method, through reverse propagation to continuously adjust weights of each layer, in an effort to minimize errors, it can structure and store massive mapping relationships hidden under a big dataset, owning excellent generalization ability and strong fault tolerance (Liu et al., 2022).
Backpropagation algorithm has two elements, in simple terms, forward propagation of input information and reverse propagation of output errors. For the former element, input parameters pass through hidden layer with weights and biases, then undergo nonlinear transformation, generating output potential. This output will be compared with the true value in dataset, then obtain loss function, the purpose of training is to minimize it. If the fitness of network cannot meet requirements, commonly indicates large loss function, we need to adjust the mapping relationship in neural network. During reverse propagation, errors will be transmitted back through hidden layer to input layer, and distributed to all neurons in each layer, then network can modify weights and biases with Gradient Descent Algorithm (Eker et al., 2023) according to errors of each neuron (Qian et al., 2023). This propagation will be repeated until network achieve desired accuracy.
@@
For the aim of improving the universality and accuracy of neural networks, dataset includes random combinations of various material parameters and solar wind plasma parameters from standard condition to extreme, such as solar wind number density changes between and , electron temperature ranging from 5eV to 20eV. Thus, neural networks can so reliably establish the mapping relationship between input parameters and output potential, making them suitable for a wide range of applications. Since the network used to calculate instantaneous potential depends heavily on the previous potential and time step, to fully consider the impact of these two parameters, we adopt 879950 pieces of data for training, each piece of data contains 12 input parameters and 1 output parameter. The specific parameters are listed in Table 1, far more than the that calculating equilibrium potential, which uses 17970 pieces of data, each piece of data contains 1- input parameters and 1 output parameter.
Before training, data normalization is considered, eliminate differences in order of magnitude between parameters, thus avoid unnecessary network errors. ReLU is used as activation function, it can compensate for the problem of Vanishing Gradient that Sigmoid and Tanh have (Kiliçarslan, 2023), helping network learn complex patterns in data.
@@ 3.2 Dynamical Multi-scale Model
We use the shape of 2016HO3 with a width-to-length ratio smaller than 0.48 to represent asteroid, regard it as a triaxial ellipsoid with semi-major axis of 29.5, 14, 14m (Li & Scheeres, 2021). As indicated in Figure 2, the simulation domain is a sphere with a radius of 500m, asteroid locates at the center of it, the Z-axis is antiparallel to the spin axis of asteroid in Section 4.1-4.3, therefore we can simulate scenarios from asteroid’s main axis being perpendicular to the solar wind to being parallel to it, while selecting either the X-axis or the Y-axis only allows for considering one of the perpendicular or parallel situations. This spin axis obliquity allows asteroid to pass through some meaningful attitudes during rotation, resulting in the greatest impact of revolution on asteroid’s surface potential, it can facilitate comparison and analysis of these special attitudes. However, this obliquity is not common in real situations. To make our simulation more suitable for realistic situations, we studied the effect of obliquity on surface potential by altering the spin obliquity in Section 4.4.
@@
This process is illustrated in Figure 3. It is worth noting that the 1D equations in Section 2 are only used for dataset construction and theoretical explanation, which help in building the neural network, accurately complete the mapping from plasma parameters and material parameters to surface potential. In our simulation, we divided the surface of the asteroid into many very small areas. Each area equals a capacitor that has the same surface potential. We use the neural network to calculate the surface potential of the capacitor instead of solving the equations in Section 2. Then we combine all the surface potential of small areas to form the 3D surface potential of the asteroid. After obtaining asteroid’s surface potential, the calculations of potential of the remaining simulation domain, electric field around asteroid, and the drift-diffusion motion of charged particles are all done in 3D.
In order to verify the reliability of our three-dimensional surface charging and particle motion model, we operate parameters of PIC simulation (Xie et al., 2023), where eV, secondary electron yield has a peak yield of at the energy of eV. Solar wind parameters are at their typical values, with number density of , a velocity of 400km/s, ions and electrons are emitted in the -X direction, time step is 1.0e-6s, while the simulation period is 0.001s. As they assume that the asteroid is static, the angle between the main axis of asteroid and plane Y-O-Z is fixed. Results are shown in Figure 4. Notably, under normal condition, to ensure a fair comparison with PIC, we set the plasma near asteroid (about 200m) to be quasi-neutral here, similar to the PIC method (Xie et al., 2023). However, this does not accurately reflect the real situation, due to the solar wind flow and electric fields in space, the potential generally does not drop to zero at 200m from asteroid (Zimmerman et al., 2014). Especially during a solar storm, the Debye length of solar wind plasma can become almost double its normal value. As a result, in simulation of Section 4.2, we will consider the plasma to be quasi-neutral 500m away from asteroid, potential around asteroid (about 50m) may still exhibit an obvious negative value.
We can see a positive potential up to +13.3V on the dayside and a negative potential as low as -37.5V on the nightside when the angle , while surface potential is about +12V near the subsolar point, and -35V on the nightside with PIC simulation.
Apart from the three situations mentioned above, we also compared the potential distribution around asteroid for Case 4 and Case 8 in reference (Xie et al., 2023). The difference between simulated parameters in Case 4 and that used for Figure 4 locates at solar wind number density, equals to in Case 4, while Case 8 differs from Figure 4’s simulated environment by . Simulated results of our method and PIC simulation with Cases4 and 8 are shown in Figure 5. Results simulated by our three-dimensional multi-scale model are on the left, and the right subfigure shows results of PIC simulation.
Our results are very close to those of previous research in terms of both curve trend and values. Given that our much larger simulation region than PIC’s, we assume that potential 500m away from asteroid is 0. Therefore, within the range -150m50m, there exists some minor differences between our method and PIC, which will not exceed 3V. It can be seen that our method is reliable.
4 Simulation
We will conduct dynamic simulations asteroids about surface charging under different conditions, the four main categories are shown as follows:
4.1 Surface Charging Properties of Rotating Asteroids under Normal Condition
@@
Figure 6 displays the surface potential of asteroid with three classical attitudes, as well as the distribution of electron density and ion density around it, the rotation period is always 1 day. Consistent with the results of static simulation, the minimum potential the asteroid can achieve on the nightside, -5.74V, is the lowest when compared with other two conditions, -4.6V when and -4.5V when . Meanwhile, if the main axis of asteroid is perpendicular to the direction of solar wind incidence, potential on the dayside is obviously higher than other situations.
Apart from period, the direction and angle of rotation can also affect charging results, for example, and are the same situation for asteroids when we use static simulation, but considering that they are rotating, their results differ significantly. Besides, from Figure 7, we can conclude that no matter how long the period is, surface potential tends to be low when solar wind is obliquely incident, this phenomenon will become more obvious as rotation period increases. That’s because when asteroid rotates to , the subsolar point , where often has the maximum potential, is close to the terminator . According to Poissons Equation, the sudden change in potential over a short distance will strengthen electric field near asteroid, accelerate electrons to m/s, the resulting higher electron density and electron temperature therefore reduce surface potential. This process continues until electric field and plasma reach a dynamic balance, which will be broken and reshaped after asteroid rotates through a large enough angle.
4.2 Impact of Rotation on Asteroids during the Passage of Solar Storm
@@
Figures 8-9 show results from a simulation with the occurrence of CME. Five attitudes are selected for display, in regard to typical solar wind, the dayside always charged positively, its potential depends on solar elevation angle, will decrease apparently during the process from subsolar point to the terminator, until it turns into a small negative potential on the nightside. When the forward shock exists, solar wind number density and temperature experience a sharp increase, solar wind starts to accelerate. At this point, a sudden drop can be seen in surface potential, photoelectron current is inferior to the current generated by high-energy electrons, the dayside is charged negatively, such as the potential of -52V when , the minimum potential on the nightside can reach -84.54V, close to 20 times that under normal circumstance, at this time, electron density here is . Therefore, even though at the same angle, potential difference between parts of asteroid’s surface can exceed 40V, contributing to exponential growth of electric field strength, as seen in Figure 9, electric field strength near asteroid, particularly the terminator, where there is a sharp transition between positive and negative potentials, has reached the level of 10V/m, greatly enhancing the sensitivity of electrons and ions to the charged properties of asteroid, plasma is extremely unstable, thus the surface potential curve performs strong volatility. It only takes 1/12 period to make the average electron density increase from to , bring about a potential drop of more than 10V.
@@
In addition to magnitude, the distribution of potential is also worth investigating. Figure 10 depicts the normalized potential, we can precisely analyze the potential distribution around asteroid, especially where the minimum potential is located. According to equation (11-12) and kinetic theory of plasma, both electrons and ions during shock achieve larger diffusion coefficients, the effect of drift-diffusion can partially overpower the flow of solar wind, increasing the density around asteroid’s backside, then decrease surface potential there, allowing electrons to gather at the terminator, where has the strongest electric field. In contrast to the virtually uniform negative potential on the nightside during typical solar wind and early CME, potential during shock has a distinct gradient. In most cases, potential at the terminator is clearly lower than other parts, except , when solar elevation angle declines most slowly on the dayside, reducing the potential difference near the terminator, thereby weakening the electric field. Because of the fast-flowing solar wind, ions produce a considerable wake near asteroids tail, whereas high-temperature electrons do not. As a result, electron current far exceeds ion current, resulting in the lowest potential here.
@@
Then we change the rotation period. Charging results of asteroids with varying periods under typical solar wind have been discussed in 4.1. We now focus on the remaining three conditions. Apart from 0.467 hour, we select three additional periods for each stage based on their respective characteristics, results are shown in Table 3, including the maximum potential , the minimum potential , the maximum electric field strength , and the range of potential differences occur on asteroids.
@@
4.3 Differences between Asteroids with Various Surface Minerals
| Plagioclase | 5.58 | 238 | 2.8 | 1000 |
| Orthopyroxene | 5.14 | 259 | 2.1 | 700 |
| Ilmenite | 4.29 | 297 | 2.5 | 600 |
Solar wind parameters take that same as 4.1, rotation periods of asteroid are 1 hour, 6 hours and 1 day.
@@ 4.3 Differences between Asteroids with Various Surface Minerals
4.4 Effect of Orbital Motion and Obliquity on Surface Charging of Rotating Asteroids
For asteroids with very long rotation periods, such as those longer than half a year, the variation in solar wind incidence angle due to orbital motion would become important. Therefore, we consider revolution based on the solar wind parameters in 4.1, plagioclase is selected as surface mineral of asteroid. Given that the orbital period of 2016HO3 is 365.9 days, we set one year as the orbital period of asteroid in the following studies, the direction of revolution is shown in Figure 12. We investigated surface charging of asteroids with rotation periods of 1 month, half a year, and 1 year under dynamic solar wind incidence directions. Results are shown in Figure 13.
Consistent with the conclusion in 4.1, the minimum potential that asteroid can reach during one rotation period still occurs when asteroid rotates to the attitude where the subsolar point is close to the terminator. For asteroid with period of 1 month, the impact of revolution is negligible. Compared with the situation where revolution is not considered in 4.1, the change in potential is within 0.05V. However, when asteroid has rotation period longer than half a year, the influence of revolution on surface charging is worth discussing. In addition to the effect on potential distribution in Figure 13, the minimum potential of asteroid’s surface during simulation has also increased, from -5.96V in 4.1 to -5.62V in 4.4 for asteroid with period of half a year, from -5.96V to -5.53V for asteroid with period of 1 year, but remained well below the minimum potential that asteroids with short periods can achieve. Considering the constantly varying solar wind incidence angle, revolution can also be seen as a disruption and restructuring to the equilibrium state around asteroid, the motion of electrons and ions near asteroid is significantly disturbed, with velocity vectors varying much more dramatically than in 4.1. This mitigates the electron accumulation phenomenon around asteroid. Thus the potential curve becomes smoother. This effect is particularly noticeable when the rotation period of asteroid is comparable to the orbital period.
Furthermore, in the above simulations, we assume that spin axis perpendicular to the X-O-Y plane, because this orientation allows asteroid to pass through some meaningful attitudes during rotation, such as those depicted in Figure 8, where and . This orientation facilitates comparison and analysis of these special attitudes during asteroid’s rotation. However, actually, asteroid’s spin axis may have an inclination relative to the orbital plane. Consequently, as illustrated in Figure 12, we will alter the spin axis to explore surface potentials of asteroids under more common conditions with obliquities of (equivalent to previous simulation), , and . Results are shown in Figure 14.
If orbital motion is ignored, obliquity has little effect on the surface charging of asteroid. Therefore, we primarily discuss the surface potential of asteroids with different obliquities taking revolution into consideration. If spin axis obliquity , asteroid’s rotational motion and orbital motion occur in two mutually perpendicular planes, resulting in the greatest impact of revolution on asteroid’s surface potential. As obliquity decreases, the potential curves gradually becomes gentler. When , because both rotation period and orbital period of asteroid are one year, making the direction of solar wind incidence stationary relative to asteroid’s surface. At this time, the maximum and minimum surface potentials of asteroid remain stable around 8.5V and -4.5V, respectively.